OpenStudy (anonymous):

Evaluate the integrals below by interpreting them in terms of areas. (a) ∫10−5|x|dx = (b) ∫90|5x−5|dx=

OpenStudy (anonymous):

\[\int\limits_{0}^{9}\left| 5x-5 \right|dx\] \[\int\limits_{-5}^{10}\left| x \right|dx\]

OpenStudy (sparrow2):